OpenStudy (anonymous):

find the sum from n=1 to infinity of (n/2^n)

OpenStudy (kainui):

http://www.wolframalpha.com/input/?i=sum%20from%20n%3D1%20to%20infinity%20of%20(n%2F2%5En)&t=crmtb01&f=rc I wish I could help more than that, but I honestly don't know.

OpenStudy (anonymous):

i tried it but the finial answer is 2 but im not sure if its right?

OpenStudy (anonymous):

Hmmmm, I wonder...

OpenStudy (anonymous):

Sometimes binomial theorem helps, but I don't think it would in this case.

OpenStudy (psymon):

Well it definitely converges.....but I'm not sure you can get an ACTUAL answer for a sum.

OpenStudy (anonymous):

i don't know which method to use. I understand how to solve sum from n=1 to infinity of (1/2^n). but because there is an n in the numerator i couldn't solve it

OpenStudy (psymon):

I don't think you can get an actual sum, I think you just prove it converges.

OpenStudy (anonymous):

the problem is the question says find the sum

OpenStudy (psymon):

Unless it's a geo-series or a telescoping series, I have no idea if there is a way to find the exact sum. Obviously there must be if it's asking for one, but that's new to me.

OpenStudy (anonymous):

I proved that its convergent using the ratio test but i dont know how to find the sum

OpenStudy (anonymous):

Hmmm, I wonder what happens if you substitute \(n\) with \(n+1\) and try to simplify...

OpenStudy (psymon):

Yeah, I did the same.

OpenStudy (psymon):

Ratio test to prove convergence and then I didn't know what more to do, lol

OpenStudy (anonymous):

\[ \frac n {2^n} \]Put in \(n+1\):\[ \frac {n+1} {2^{n+1}} = \frac1 2\left(\frac {n+1}{2^n}\right) = \frac 1 2 \left( \frac n {2^n} + \frac 1 {2^n}\right) \]

OpenStudy (anonymous):

I dunno if this helps but it's something to think about.

OpenStudy (anonymous):

\[ a_n = \frac 1 2 \left(a_{n-1} + \frac 1 {2^n} \right) \]

OpenStudy (tkhunny):

Old trick 1) \(\dfrac{1}{2} + \dfrac{2}{2^{2}} + \dfrac{3}{2^{3}} + \dfrac{4}{2^{4}} + ... = S\) Multiply by 1/2 (Doesn't seem like it will get us anywhere, does it?) 2) \(\dfrac{1}{2^{2}} + \dfrac{2}{2^{3}} + \dfrac{3}{2^{4}} + \dfrac{4}{2^{5}} + ... = \dfrac{S}{2}\) Subtract 2 from 1 \(\dfrac{1}{2} + \dfrac{1}{2^{2}} + \dfrac{1}{2^{3}} + \dfrac{1}{2^{4}} + ... = \dfrac{S}{2}\) Now, the left hand side is just a geometric series. \(\dfrac{\dfrac{1}{2}}{1 - \dfrac{1}{2}} = 1 = \dfrac{S}{2} \implies S = 2\)

OpenStudy (anonymous):

@tkhunny The step in which you subtracted 2 from 1, what exactly did you do?

OpenStudy (psymon):

Yeah, thats what Im looking at.

OpenStudy (tkhunny):

It sort of telescopes. Each denominator matches up nicely between 1) and 2), with the exception of the first term.

OpenStudy (anonymous):

this is what i did so far \[\sum_{1}^{\infty} (\frac{ n }{ 2^n }) = \sum_{1}^{\infty} (\frac{ n-1+1 }{ 2^n }) = \sum_{1}^{\infty} (\frac{ n-1 }{ 2^n } +\frac{ 1 }{ 2^n } )= \sum_{1}^{\infty} (\frac{ n-1 }{ 2^n }) + \sum_{1}^{\infty} (\frac{ 1 }{ 2^n } )\] and I know that \[\sum_{1}^{\infty} (\frac{ 1 }{ 2^n }) = 1\] which means that \[\sum_{1}^{\infty} (\frac{ n-1 }{ 2^n }) =1\] but how to prove that \[\sum_{1}^{\infty} (\frac{ n-1 }{ 2^n })\] is really equals 1

OpenStudy (anonymous):

@Brittny You are basically circling the problem. To prove that the n-1/2^n equals one, we really need to prove that n/2^n equals 2 which brings us back to where we started.

OpenStudy (anonymous):

@tkhunny I see what you did now. I'm trying to generate another solution though. Let's hope it works out =] Btw that's the kind of trick one uses to prove the sum formula for an arithmetic sequence; but it doesn't always come to mind ;]

OpenStudy (anonymous):

Then you do the subtraction.

OpenStudy (tkhunny):

Standard fare for linearly-increasing benefits from the world of actuaries. :-)

OpenStudy (anonymous):

@tkhunny u becoming an actuary?

OpenStudy (anonymous):

@tkhunny why did u multiply by 1/2 it seems that any number could work and i would get an infinite number of solutions

OpenStudy (tkhunny):

@genius12 ahaha! No, already been doing that for 30 years! @Brittny 1/2 is the common ratio for what I knew ahead of time would be a geometric series.. Always the common ratio.

OpenStudy (anonymous):

@tkhunny WAAAAAAAAAAAAAAAAAT?@!?#!@?#@!#?!@#@?!#?# SRS BRO? lol i thought u were my age =.=

OpenStudy (tkhunny):

That picture of me singing opera? It's more than 15 years old!

OpenStudy (tkhunny):

Sorry, wrong opera. That one is only about 5 years old. My bad.

OpenStudy (anonymous):

to find the sum of the seires. we took in school its done by two ways , making the seires as geo seires or telescoping series

OpenStudy (anonymous):

i don't know if that helps

OpenStudy (psymon):

Wish there was another example for that trick.

OpenStudy (anonymous):

I think this is how you would generate a formula for these type of series. First start with a finite geometric series:\[1+x+x^2+\cdots +x^n=\frac{x^{n+1}-1}{x-1}.\]Then take the derivative of both sides:\[1+2x+3x^2+\cdots +nx^{n-1}=\frac{(n+1)x^n(x-1)-(x^{n+1}-1)}{(x-1)^2}\]\[=\frac{nx^{n+1}-(n+1)x^n+1}{(x-1)^2}.\]Multiply both sides by x:\[x+2x^2+3x^3+\cdots +nx^n=\frac{nx^{n+2}-(n+1)x^{n+1}+x}{(x-1)^2}\]In summation form this is saying:\[\sum_{k=1}^{n}kx^k=\frac{nx^{n+2}-(n+1)x^{n+1}+x}{(x-1)^2}\]Now take n to infinity. If |x|<1, it converges, since both:\[nx^{n+2},(n+1)x^{n+1}\]go to zero. If |x|<1, we end up with:\[\sum_{k=0}^{\infty}kx^k=\frac{x}{(x-1)^2}.\]If you let x = 1/2, then you get 2 as an answer. Not entirely sure if this is correct, I'll try it out with some other series of this form.

OpenStudy (anonymous):

It works for:\[\sum_{n=1}^\infty \frac{n}{3^n}=\frac{3}{4}\] http://www.wolframalpha.com/input/?i=sum+from+n%3D1+to+infinity+of+%28n%2F3%5En%29 and:\[\sum_{n=1}^\infty \frac{n}{5^n}=\frac{5}{16}\] http://www.wolframalpha.com/input/?i=sum+from+n%3D1+to+infinity+of+%28n%2F5%5En%29

OpenStudy (anonymous):

thanks a lot

OpenStudy (tkhunny):

@Brittny "...find the sum of the seires. We took in school its done by two ways , making the series as geo series or telescoping series " This is good. In this case, in my demonstration, the original series was made to "telescope" into a geometric series. The is a little different from telescoping that annihilates everything, but it is still the familiar "telescoping". You need both concepts to solve this one.